# TabRep: a Simple and Effective Continuous Representation for Training Tabular Diffusion Models

### v2

@@ 3.1 Tabular Diffusion Framework
Unified Continuous Diffusion. Training diffusion models on a separate continuous-discrete data representation comes with the challenge of joint optimization. An alternative is to model a unified continuous data representation that circumvents these challenges. [24] introduces StaSy which utilizes a naive one-hot encoding representation to unify the data space but necessitates additional computationally heavy tricks to improve its performance. StaSy trains a diffusion model on a subset of data before gradually extending to the whole training set. Thus, in addition to the diffusion model parameter, $\theta$, the model is required to learn a selection importance vector, $\mathbf{v}\in[0,1]$ through an alternative convex search [2]. The search solves the loss function by iteratively optimizing between $\theta$ and $\mathbf{v}$ while also requiring a self-paced regularizer, $r(.)$, that requires fine-tuning. Another method by [58] introduces TabSYN that unifies the data space by transforming discrete features using a learned $\beta$-VAE latent embedding. The architecture of the VAE consists of a tokenizer that encodes discrete features into a one-hot vector, then parameterizes each category as a learnable vector learned by a Transformer [52] encoder. Likewise to StaSy, TabSYN is also computationally expensive since it relies on training a separate VAE before the diffusion process to obtain latent codes. Additionally, the quality of synthetic data is highly dependent on the quality of the latent space. In the following section, we show that leveraging our simple and effective continuous data representation to train tabular diffusion models is essential in synthesizing high-quality tabular data.
@@ 4.1 Geometric Implications on the Data Representation
Diffusion models rely on gradients derived from a learned vector field to denoise data iteratively [20, 31]. For regions within proximity of the singular hyperplanes, gradients suffer from high variance due to conflicting directions arising from equidistant one-hot points. We show that the variance of the score function increases asymptotically with the degree of $n$-singular points. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as: $p_{t}(x|e_{k})=\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)$. We derive the variance of the conditional score function evaluated at an $n$-singular point as:
@@
where we define the conditional score and expected score as $g$ and $\bar{g}$. See Appendix A.1 for a proof. We find that near an $n$-singular point, $x$, the posterior-weighted variance of the conditional score function is strictly positive and increases asymptotically with $n$. In contrast, at a non-singular point, the posterior leans towards $e_{k^{*}}$, leading the score variance to approach zero.
@@ 4.2 TabRep Architecture
In addition to nominal features, the separability of CatConverter coupled with its circular geometry naturally accommodates both periodic and ordinal features. This facilitates an enhanced preservation of the feature’s intrinsic nature and characteristics. Note that prior to CatConverter, we introduce a static one-dimensional embedding, Dictionary (dic), but we find that it underperforms.
@@ A.1 Variance of Learning Gradients in Diffusion Models
We show that the variance of the score function at an $n$-singular point increases asymptotically with respect to $n$ dimensions. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as:
@@
$p(e_{k})=\frac{1}{K}\quad\forall k\in\{1,\ldots,K\}$ (20)
@@
And $\forall m\notin S$, $d_{m}(x)^{2}>d_{k}(x)^{2}$, thus:
@@
We now compute the expected score: $x_{t}=\mathcal{N}(\alpha_{t}x_{0},b_{t}^{2})$
@@
We find that at an $n$-singular point $x$, the posterior-weighted variance of the score function is strictly positive and increases asymptotically with $n$. In contrast, at a non-singular point, the posterior leans towards $e_{k^{*}}$:

### v3

@@ 3.1 Tabular Diffusion Framework
Unified Continuous Diffusion. Training diffusion models on a separate continuous-discrete data representation comes with the challenge of joint optimization. An alternative is to model a unified continuous data representation that circumvents these challenges. StaSy [24] utilizes a naive one-hot encoding representation to unify the data space but necessitates additional computationally heavy tricks to improve its performance. It trains a diffusion model on a subset of data before gradually extending to the whole training set. Thus, in addition to the diffusion model parameter, $\theta$, the model is required to learn a selection importance vector, $\mathbf{v}\in[0,1]$ through an alternative convex search [2]. The search solves the loss function by iteratively optimizing between $\theta$ and $\mathbf{v}$ while also requiring a self-paced regularizer, $r(.)$, that requires fine-tuning. Another tabular diffusion method, TabSYN [58], unifies the data space by transforming discrete features using a learned $\beta$-VAE latent embedding. The architecture of the VAE consists of a tokenizer that encodes discrete features into a one-hot vector, then parameterizes each category as a learnable vector learned by a Transformer [52] encoder. Likewise to StaSy, TabSYN is also computationally expensive since it relies on training a separate VAE before the diffusion process to obtain latent codes. Additionally, the quality of synthetic data is highly dependent on the quality of the latent space. In the following section, we show that leveraging our simple and effective continuous data representation to train tabular diffusion models is essential in synthesizing high-quality tabular data.
@@ 4.1 Geometric Implications on the Data Representation
Diffusion models rely on gradients derived from a learned vector field to denoise data iteratively [20, 31]. For regions within proximity of the singular hyperplanes, learning the gradients of diffusion models suffers from high variance due to conflicting directions arising from equidistant one-hot points. We show that the variance of the conditional score function increases asymptotically with the degree of $n$-singular points. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as: $p_{t}(x|e_{k})=\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)$. We derive the variance of the conditional score function evaluated at a minimal $n$-singular point as:
@@
where we define the conditional score and expected score as $g$ and $\bar{g}$. See Appendix A.1 for a proof. We find that near a minimal $n$-singular point, $x$, the posterior-weighted variance of the conditional score function is strictly positive and increases asymptotically with $n$. In contrast, at a non-singular point, the posterior leans towards $e_{k^{*}}$, leading the score variance to approach zero.
@@ 4.2 TabRep Architecture
In addition to nominal features, the separability of CatConverter coupled with its circular geometry naturally accommodates both periodic and ordinal features. This facilitates an enhanced preservation of the feature’s intrinsic nature and characteristics. Note that prior to CatConverter, we introduced a static one-dimensional embedding, Dictionary (dic), but we found that it underperformed.
@@ A.1 Variance of Learning Gradients in Diffusion Models
We show that the variance of the score function at a minimal $n$-singular point increases asymptotically with respect to $n$ dimensions. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as:
@@
$p(e_{k})=\frac{1}{K}\quad\forall k\in\mathcal{S}=\{1,\ldots,K\}$ (20)
@@
where $d$ is defined in Equation 7. $\forall m\notin S$, $d_{m}(x)^{2}>d_{k}(x)^{2}$, thus:
@@
We now compute the expected score:
@@
We find that at a minimal $n$-singular point $x$, the posterior-weighted variance of the score function is strictly positive and increases asymptotically with $n$. In contrast, at a non-singular point, the posterior leans towards $e_{k^{*}}$:
@@
We now generalize the above result to an arbitrary categorical prior $\Pi=\{\pi_{k}\}_{k\in\mathcal{S}}$ over categories $\{e_{k}\}_{k\in\mathcal{S}}$, where $x_{0}\sim\Pi$ and the forward diffusion process is defined as:
$p_{t}(x|x_{0})=\mathcal{N}(x|\alpha_{t}x_{0},\sigma_{t}^{2}I).$ (48)
The marginal likelihood is then given by:
$p_{t}(x)=\sum_{k\in\mathcal{S}}p_{t}(x|x_{0}=e_{k})\pi_{k}=\sum_{k\in\mathcal{S}}\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)\pi_{k}.$ (49)
Suppose we observe a noised sample $x\in\mathbb{R}^{k}$ at time $t$. Then the posterior probability that $x$ was generated by adding noise to $e_{k}$ is:
$q_{t}(x_{0}=e_{k}|x)=\frac{p_{t}(x|x_{0}=e_{k})\pi_{k}}{\sum_{m\in\mathcal{S}}p_{t}(x|x_{0}=e_{j})\pi_{m}}.$ (50)
We compute the expected score at time $t$ as follows:
$\mathbb{E}_{p_{t}(x)}[\nabla_{x}\log p_{t}(x)]$ $=\mathbb{E}_{x_{0}\sim\Pi}\left[\mathbb{E}_{p_{t}(x|x_{0})}\left[\nabla_{x}\log p_{t}(x|x_{0})\right]\right]$ (51)
$=\sum_{k\in\mathcal{S}}\pi_{k}\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)\nabla_{x}\log\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I).$ (52)
The score of the Gaussian is:
$\nabla_{x}\log\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)=-\frac{1}{\sigma_{t}^{2}}(x-\alpha_{t}e_{k})=\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}.$ (53)
Therefore, the expected score becomes:
$\bar{g}:=\mathbb{E}_{p_{t}(x)}[\nabla_{x}\log p_{t}(x)]$ $=\sum_{k\in\mathcal{S}}\pi_{k}\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)\cdot\left(\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}\right)$ (54)
$=C\sum_{k\in\mathcal{S}}\pi_{k}\exp\left(-\frac{\|x-\alpha_{t}e_{k}\|^{2}}{2\sigma_{t}^{2}}\right)\left(\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}\right),$ (55)
where $C:=\frac{1}{(2\pi)^{k/2}\sigma_{t}^{k}}$.
Next, we compute the variance of the conditional score around its expectation:
$\mathbb{E}_{x_{0}\sim\Pi}\left[\mathbb{E}_{p_{t}(x|x_{0})}\left[\left\|\nabla_{x}\log p_{t}(x|x_{0})-\bar{g}\right\|^{2}\right]\right]$ $=\sum_{k\in\mathcal{S}}\pi_{k}\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)\left\|\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}-\bar{g}\right\|^{2}$ (56)
$=C\sum_{k\in\mathcal{S}}\pi_{k}\exp\left(-\frac{\|x-\alpha_{t}e_{k}\|^{2}}{2\sigma_{t}^{2}}\right)\left\|\frac{\alpha_{t}e_{k}-x}{\sigma_{t}^{2}}-\bar{g}\right\|^{2}.$ (57)
This formulation reveals that score variance is low when the posterior is sharply peaked (i.e., $x$ is close to a single one-hot vector), and increases when the posterior mass is spread over multiple categories. The minimal $n$-singular point case is recovered when $\pi_{k}=\frac{1}{n}$ for $k\in\mathcal{S}$ and $x=\frac{1}{n}\sum_{k\in\mathcal{S}}e_{k}$, confirming the consistency of both analyses.